OpenStudy (anonymous):

The answer is 10/3 mph, can someone help me with the equation? Sam can row 10 miles downstream in one third the time he can row 6 miles upstream. If Sam can row 5 miles per hour in still water, how fast is the current?

OpenStudy (amistre64):



OpenStudy (amistre64):

lets time him :) in 1hour he can go 6 miles up; and in 1/3 hour he can go 10 miles down or 10 down in and hour and 6 up in 3

OpenStudy (anonymous):

s = d/t (=) t = s /v so if you equalize the times, 10/(5+x) = 6/(3(5-x)) blablabla... 5+x=25-5x 20 = 6x x = 10/3

OpenStudy (anonymous):

t = d / s i meant

OpenStudy (anonymous):

and ofc (5+x) is he swimming 5mph plus the current (x) and (5-x) is he swimming minus the current force

OpenStudy (anonymous):

ok.. Step by step: 10 6 5+x 5-x 3 5+x 5-x ---- = ----- (=) ----- = ---- x -- (=) ----- = ----- (=) 5+x = 5(5-x) 5+x 3(5-x) 10 1 6 10 2 5+x = 5(5-x) (=) 5+x=25-5x (=) 20 = 6x (=) x = 10/3

OpenStudy (anonymous):

awesome. Thank you much